Please help me answer this.
The problem asks "Find the difference quotient and simplify your answer."
f(x)= x^3 + x, f(x+h)-f(x)/h, h=0

Answer by stanbon(75887)

"Find the difference quotient and simplify your answer."
f(x)= x^3 + x, [f(x+h)-f(x)]/h
-------
f(x+h)= (x+h)^3+x
f(x+h)= x^3+3x^2h+3xh^2+h^3 + (x+h)
f(x) = x^3 + x
-----
Subtract f(x) from f(x+h) to get:
f(x+h)-f(x) = 3x^2h + 3xh^2 + h^3 + h
-----
Divide by "h":
[f(x+h)-f(x)]/h = 3x^2 + 3xh + h^2 + 1
============================================
Cheers,
Stan H.
==================